### v1

# Exploring the AGN Fraction of a Sample of JWST’s Little Red Dots at : Overmassive Black Holes Are Strongly Favored

Abstract
JWST is revolutionizing our view of the early Universe by pushing the boundaries of detectable galaxies and black holes in redshift (upward) and mass (downward). The Little Red Dots (LRDs), detected by several surveys at , present a significant interpretational challenge, as their spectral energy distributions can mimic both AGN and stellar population templates. This study analyzes eight LRDs from the CEERS survey, utilizing NIRCam and MIRI photometry. By performing SED fitting across a vast parameter space, we explore a broad range of AGN fractions, defined as the ratio of the monochromatic luminosities (AGN, galaxy, and dust) at 0.53 rest-frame. We find that the SEDs of all the LRDs investigated are consistent with having significant AGN contributions, with the best-fitting AGN fractions ranging between 40% and 85%. Moreover, assuming these LRDs do indeed host AGN, we can place limits on their black hole masses using the inferred AGN bolometric luminosities and adopting the Eddington limit. We find that, independent of the specific AGN fraction adopted, the LRDs’ black holes are significantly overmassive relative to their host galaxies—by dex, and up to dex in the most extreme cases—compared to the local relation. The presence of overmassive black holes in the high- Universe may provide the strongest evidence yet of heavy black hole seeding occurring during the cosmic dark ages.
1 Introduction
JWST provides an unprecedented view of the early Universe. Unlike earlier telescopes that observed massive quasars at (Fan et al., 2001, 2003, 2006; Cool et al., 2006; Willott et al., 2009; Wang et al., 2007; Willott, 2011; Mortlock et al., 2011; Wu et al., 2015; Bañados et al., 2018; Fan et al., 2023), where the total spectral energy distribution (SED) is dominated by the emission of the central supermassive black hole (SMBH, with mass ), JWST can often capture the light from the host’s stellar component (see, e.g., Ding et al. 2023), especially in the case of lower-mass (i.e., ), lower-luminosity black holes (i.e., ). This extraordinary sensitivity has led to the detection of new populations of SMBHs in the high- Universe, farther and with lighter and more compact hosts (e.g., GN-z11 Maiolino et al. 2024a; Baggen et al. 2023; Greene et al. 2024).
Observations of high- systems play an essential role in our ability to understand the early evolution and co-evolution of black holes and their host galaxies. Extensive studies in the local Universe (i.e., ) have shown the existence of evolutionary relationships between the central SMBHs and their hosts, at least in massive galaxies (see, e.g., Ferrarese & Merritt 2000; Gebhardt et al. 2000; Kormendy & Ho 2013). These observations led to empirical relations between the mass of the central black hole and several properties of the host. For example, the with the stellar mass of the host galaxy, the with its velocity dispersion, and the with the dynamical mass of the host (Hu, 2008; Kormendy & Ho, 2013; Heckman & Best, 2014; Reines & Volonteri, 2015). While nuances in how these parameters are measured could have some underlying effect on the normalization and scatter in the respective relationships (e.g., Shankar et al. 2019), the overall trend shows an undeniable link between the growth of SMBHs and their host galaxies, at least in the local Universe. This co-evolution is not yet well investigated in the high- Universe, primarily because the detection of the stellar component of galaxies hosting very massive black holes was rare in the pre-JWST era.
Efforts to quantify these scaling relations using JWST observations at have revealed a population of central SMBHs that are overmassive compared to the stellar content of their hosts (Maiolino et al., 2024a, 2023; Harikane et al., 2023; Larson et al., 2023; Übler et al., 2023; Stone et al., 2023; Furtak et al., 2024; Kocevski et al., 2023; Kokorev et al., 2023; Yue et al., 2024b; Juodžbalis et al., 2024; Kocevski et al., 2024), but seem to agree with the stellar velocity dispersion and the dynamical mass (Maiolino et al., 2023). More specifically, the relation inferred from JWST detections deviates by from the local relationship: black holes are overmassive by factors of to with respect to local counterparts hosted by a galaxy of the same size (Pacucci et al., 2023).
The prevalence of such overmassive black holes brings into question the seeding mechanism responsible for the formation of such high- systems (Pacucci et al., 2023; Volonteri et al., 2023; Natarajan et al., 2024; Pacucci & Loeb, 2024). Heavy seed scenarios (see, e.g., Volonteri et al. 2008, and the reviews by Woods et al. 2019; Inayoshi et al. 2020) were first introduced to explain the extremely massive SMBHs observed at (see, e.g., Mortlock et al. 2011; Wu et al. 2015; Bañados et al. 2018), which would otherwise lack the necessary time to growth from standard, low-mass seeds. These heavy seeding models refer to various channels that may have shaped the formation of black holes in the early Universe, such as direct collapse black holes (Loeb & Rasio, 1994; Bromm & Loeb, 2003; Lodato & Natarajan, 2006). Some recent studies (e.g., Pacucci et al. 2023; Natarajan et al. 2024; Scoggins et al. 2023; Pacucci & Loeb 2024) suggest that the high- overmassive black holes observed by JWST may be the direct result of this heavy seeding scenario.
However, determining the location of these high- systems in the plane crucially depends on how the contributions from the stellar and the active galactic nucleus (AGN) components of the SEDs are evaluated, as well as the individual uncertainties in the determination of the stellar mass and the black hole mass. This study investigates how different assumptions on the AGN/star ratio (i.e., how much these two components generate the observed flux) affect the goodness-of-fit with the observed photometry and the system’s location on the plane.
In particular, we analyze eight extremely red objects (EROs) first reported in Barro et al. 2024 and observed with JWST. As detailed in Sec. 2, these EROs are a subset of the more general population of “Little Red Dots”, or LRDs, a name that was first introduced by Matthee et al. (2023) and that is now of widespread use. Since the dataset from Barro et al. (2024) was published, several other studies reported the detection of similarly red and compact objects, which are ubiquitous in many JWST fields (see, e.g., the extensive catalog provided by Kocevski et al. 2024). In particular, Greene et al. (2024) argued that most (i.e., ) of these sources host a SMBH at their center, based on the detection of broad line regions (BLRs).
The rest of this paper is organized as follows: Section 2 discusses the data used and summarizes the data reduction process. Section 3 details our SED modeling and choices of parameters. In Section 4 and 5, we discuss results from the SED modeling and its implication within the larger context of galaxy and black hole co-evolution.
2 Data
This study is based on the sample of EROs by Barro et al. (2024), which was the first study to analyze these remarkably red and compact sources thoroughly.
Barro et al. (2024) report the discovery of 37 EROs, whose selection is based on NIRCam colors F277W-F444W 1.5 mag. These objects are selected from the JWST’s Cosmic Evolution Early Release Science Survey (CEERS, Finkelstein et al. 2023), which surveys arcmin2 of the Extended Groth Strip (EGS). Out of the 37 sources in Barro et al. (2024), we select a sub-sample of eight sources, which are characterized by spectra acquired with NIRSpec, within and , and MIRI, within and .
These objects were observed as very red galaxies in the JWST/NIRCam filters and additionally detected in either the NIRSpec filters or the MIRI filters. A common feature in all eight galaxies is a bump in the redder part of the photometric spectrum, often attributed to AGN emission.
2.1 A Note on Terminology: Extremely Red Objects vs. Little Red Dots
The name “Extremely Red Objects”, or EROs, precedes that of “Little Red Dots”, or LRDs, which has now become a standard nomenclature for the compact and strikingly red objects commonly found by JWST between and (Kocevski et al., 2024).
The color selection performed by Barro et al. (2024) for the EROs is consistent with the current definition of LRDs (see, e.g., the extensive catalog provided in Kocevski et al. 2024); however, the selection used in Barro et al. (2024) was slightly different. This paper selected extremely red sources with F277W-F444W , but no selection on morphology or rest-frame UV color was performed. Remarkably, the vast majority of sources with the F277W-F444W color selection are also point-like and have blue F150W-F200W colors, which fit the current definition of LRDs (Labbe et al., 2023; Kokorev et al., 2023; Kocevski et al., 2023, 2024). However, the subset studied in this paper do not have rest frame optical spectroscopy; most objects in the general population of LRDs have broad line emission consistent with Type 1 (unobscured) AGN (Greene et al., 2024). Additionally, and as discussed below, the population studied here are best modeled as Type 2 (obscured) AGN, suggesting they may represent a particular subset of LRDs.
3 Methods: SED modeling
Module | Parameter | Symbol | Values |
Star Formation History | e-folding time (main) | 0.1, 0.4, 0.8, 1, 2, 3, 4, 5 Gyrs | |
Main stellar age | 0.1, 0.2, 0.4, 0.5, 0.6, 0.7 Gyrs | ||
e-folding time (burst) | 0.01, 0.05, 0.1 Gyrs | ||
sfhdelayed | Age of the burst population | 0.01, 0.03, 0.05 Gyrs | |
Burst fraction | 0, 0.001, 0.01, 0.1, 0.2, 0.3, 0.4, 0.5 Gyrs | ||
Simple stellar population | Initial mass function | … | Chabrier |
bc03 | Mettalicity | 0.02 | |
Nebular Emission | Ionization parameter | log | -2.0 |
nebular | Gas Metallicity | 0.02 | |
AGN emission | Optical depth at 9.7 | 7 | |
Opening angle | 40° | ||
Viewing angle | 30° | ||
Skirtor 2016 | AGN contribution | fracAGN | 0.1 to 0.99 |
Wavelength range for fracAGN | 0.53 m | ||
Extinction in polar direction | E(B-V) | 1, 2, 3, 4, 5, 6 |
Our analysis uses the CIGALE (Burgarella et al., 2005; Boquien et al., 2019) SED fitting code. CIGALE models the far-UV to radio spectrum of galaxies to estimate physical properties such as stellar mass, AGN luminosity, star formation rate, and dust attenuation. CIGALE composite models are built based on templates describing various aspects of the galactic emission, such as the star formation history (SFH), nebular emissions, dust emission and attenuation, and the AGN contribution.
We fit the available photometry of the eight LRDs using SEDs built on a combination of two emission templates: stellar and AGN. We selected templates from the ones provided in CIGALE to vary only a small subset of parameters to focus our analysis on how the AGN fraction affects the SED fitting results. Table 1 lists the selected CIGALE templates and the parameters that were kept free, alongside the parameters where we adopted a different value from the default value. We note that we adopted the fiducial values of redshift for each source, as provided by Barro et al. (2024); hence, CIGALE does not fit redshifts.
For the star formation history, we use a standard delayed model with exponential burst: sfhdelayed. We allow the -folding time to vary from 0.1 to 5 Gyrs and the main stellar age from 0.1 Gyrs to the age of the Universe at the observed . These values are typically used in the literature for similar analyses. As these galaxies could be undergoing rapid evolution in their stellar population, we allow a range of the e-folding time for the burst population (0.01 to 0.1 Gyr), the age of the burst population (0.001, 0.05 Gyr), and the burst fraction (0 to 0.5).
For the simple stellar population, we use the bc03 module (Bruzual & Charlot, 2003), with a Chabrier (2003) IMF. Although the module defaults to solar metallicity ( = 0.02), we find that changing this parameter had little effect on our modeling. We included nebular emissions using the nebular module Bruzual & Charlot (2003). For dust attenuation, we employ the dustatt_modified_starburst module based on Charlot & Fall (2000). We employ the dl2014 module Draine et al. (2014) for galactic dust emission. For spatially unresolved sources at high redshift, there is little information regarding the relative contributions of the torus and the interstellar medium to the total dust attenuation (Hickox & Alexander, 2018). For our analysis, we do not vary dust attenuation on a galactic scale, instead choosing to focus on the torus for simplicity.
For the AGN component, we employ the skirtor2016 module based on clumpy AGN models from Stalevski et al. (2012) and Stalevski et al. (2016). We fix the parameter (i.e., the AGN optical depth at 9.7 m) to 7 and the opening angle of the torus to 40°. As skirtor uses the inclination to determine the AGN type, we allow this parameter to vary between 30°(type II) and 70°(type I)111 Note that CIGALE uses a convention where type II AGN is defined as [0°, 90°-], and type I AGN is defined as [90°-, 90°], where is the opening angle. This definition is inverted with respect to normal conventions.
The AGN fraction in skirtor is defined as the ratio between the AGN luminosity and the total luminosity over a wavelength range defined by the user. We set the (rest-frame) wavelength to 0.53 m for our analysis. We then fix the AGN fraction between 5% and 99% to determine the best set of parameters at a given AGN fraction. We changed the torus dust attenuation parameter from its default value of 0.003 and allowed it to vary between 1 and 6, as we expect this range of E(B-V) values to better match the level of obscuration reported for JWST galaxies at .
4 Results
In this Section, we first investigate how a range of AGN fraction values affects the goodness-of-fit descriptors. Then, we study how the inferred best-fit AGN fraction localizes the source in the plane.
4.1 SED models
During the SED modeling process, we found specific phenomenological trends that persisted in all analyzed LRDs. The residuals of the CIGALE models, as in the example of Fig. 1, are typically noisier at longer wavelengths. We highlight two possible explanations for this effect. First, the nebular emissions in CIGALE seem not to be well accounted for in the composite spectrum. Despite our efforts, the nebular peaks in Fig. 1 (at ) consistently show the largest residuals from the model. Second, the longer IR wavelengths are covered by MIRI, which has the shallowest coverage for the set of photometry we are modeling. In addition to larger residuals, we find that, for the LRDs that lack MIRI coverage, CIGALE modeling is biased towards an IR luminous AGN template. This strongly suggests that deep MIRI coverage is needed to accurately constrain the AGN contribution in the SEDs of LRDs (Greene et al., 2024).
Our initial approach was to test the notion that LRDs are high- obscured AGNs by fitting our sample objects as type I (unobscured) and type II (dust-obscured) AGNs. However, we found that type I AGN models failed to produce the correct emission profile to match the IR bump in our SEDs, resulting in significantly worse reduced when compared to type II models. Therefore, we decided to fit all our LRDs as Type II AGNs.
4.2 Stellar-Dominated or AGN-Dominated?
This analysis aims to constrain the nature of the LRDs while accounting for our lack of information on precisely how much the AGN contributes to the observed SED. To this end, we perform a complete analysis of how a varying AGN fraction modifies the location of the galaxy in the plane; we then use the reduced statistic provided by CIGALE to determine the best-fit parameters.
The direct outputs from the SED fitting process outlined in Section 3 are the stellar mass of the host and the AGN bolometric luminosity. This latter parameter is then used to derive the mass of the central black hole by assuming that accretion occurs at the Eddington luminosity. Note that most of the lower-luminosity broad line AGN found at by JWST thus far have a high mean Eddington ratio of close to unity (see the related discussion in, e.g., Harikane et al. 2023; Maiolino et al. 2023; Pacucci & Loeb 2024). It is important to remark that lower values of the Eddington ratios would require a larger black hole mass, which would render our black holes even more overmassive. In principle, super-Eddington accretion rates may be possible in these high- environments given the wide availability of cold gas (Power et al., 2010; Pacucci & Loeb, 2020). However, for the purposes of this analysis, we assume a maximum luminosity of approximately , as has been observed in other AGN populations across a range of redshifts (e.g., Aird et al. 2012; Fan et al. 2023; Li et al. 2024).
This process provided estimates for the black hole and stellar masses for any given AGN fraction. An example of this study of the parameter space is shown in Figure 2; as the AGN fraction increases (i.e., higher AGN contribution to the overall SED), so does the inferred black hole mass, which then corresponds to a decrease in stellar mass. This observed behavior matches our expectations based on how CIGALE uses composite models to construct SEDs.
The sharp drop that can be seen at an AGN fraction of is likely due to the absence of MIRI coverage at longer wavelengths, leading to a preference for a highly reddened AGN component. By considering models where the AGN is highly reddened, CIGALE no longer varies the AGN contribution in response to a change in AGN fraction. Rather, the flux of the galaxy component is lowered to compensate for the additional contribution to the optical luminosity from the IR-luminous AGN component.
For the eight sources analyzed, the AGN fractions yielding the lowest chi-squared values range between 40% and 85%, suggesting that the SEDs of these systems are consistent with a strong AGN contribution. However, the precise value of the AGN fraction is not well constrained; further analysis, such as spectroscopic follow-up, would be required to confirm the contribution of the AGN to the total luminosity in these sources.
4.3 Location in the Plane
The position of the eight LRDs on the plane for their best-fit models are shown by stars in Fig. 3. Generally, we find black hole masses ranging from to .
We have also plotted the Reines & Volonteri (2015) relation and the inferred high relation from Pacucci et al. (2023). Except for NC5-9553, we find that the black holes in these galaxies are overmassive by orders of magnitude (and up to in extreme cases) compared to their local Universe counterparts. Hence, these specific LRDs are well described by a scaling relation in the plane that favors overmassive black holes, such as the one presented in Pacucci et al. (2023).
Furthermore, we highlight that the presence of a black hole in these galaxies is supported by considerations on the limited amount of stellar mass that these hosts can contain, as they are very compact in size (see, e.g., Pacucci & Loeb 2024). In fact, typical effective radii are in the order of pc (Baggen et al., 2023). SED fitting executed with only stellar components (i.e., an AGN fraction of zero) would lead, in some cases, to average stellar densities that are larger, by factors , than core densities in globular clusters (Pacucci et al., in prep.).
The same figure displays the track in the plane that is due to a varying AGN fraction (in the range 5% to 99%) for each LRD. The tracks show that even assuming a modest AGN contribution ( 10%), SED fitting strongly suggests the presence of an overmassive black hole. Furthermore, for all eight LRDs, the tracks evolve to the plot’s left side, showing that as the AGN fraction increases, the objects become more overmassive on the plane. We argue that this result shows that if the LRDs observed by JWST are AGN hosts, they must be overmassive regardless of how much the AGN contributes to the overall SED.
5 Discussion and Conclusions
One of JWST’s most surprising findings to date is the omnipresence of extremely red, compact objects in the high- Universe (Kocevski et al., 2024). These faint, remarkably red sources have been selected in several surveys of different regions of the sky.
However, selecting these objects is challenging due to the degeneracy of the many processes that can cause a galaxy to look red. JWST photometry can look redder due to high equivalent width emission lines, which boosts the photometry in the JWST NIRCAM filters. Obscured AGNs could also contaminate the selection because the red power-law-like emission could result in red optical-to-infrared colors. Additionally, some LRDs have been re-classified as brown dwarfs (Nonino et al., 2023; Hainline et al., 2024; Kocevski et al., 2024).
Hence, selecting high- galaxies based only on their red colors is associated with significant uncertainties. In addition, obtaining physical insights from fitting photometric SEDs can be challenging, as they can be modeled with components that are star-dominated or AGN-dominated while maintaining reasonable goodness-of-fit.
However, despite these intrinsic limitations, our work has shown that rigorously probing the parameter space in the stellar and black hole mass can constrain our sample of LRDs as high- AGN hosts, as also suggested by previous studies (Greene et al., 2024) of different samples. In our sample of eight LRDs, we found that the best fit AGN contribution is between 40% to 85%. The inferred AGN bolometric luminosities show black holes that are significantly overmassive compared to the stellar mass of the host, based on expectations from the local Reines & Volonteri (2015) . These sources are in better agreement with the high-z relation inferred by Pacucci et al. (2023) from JWST galaxies in the redshift range .
We also found that JWST NIRCam photometry without MIRI coverage is not sensitive enough to properly constrain the AGN IR luminosity, resulting in excessively IR-luminous AGNs. Hence, based on our findings, we argue that deep MIRI coverage is essential to reveal the highly obscured population of the black hole population in the LRDs.
The fact that most of the black holes observed thus far by JWST at are overmassive with respect to the host’s stellar content is not necessarily surprising. Earlier works (e.g., Volonteri et al. 2023) predicted that if JWST were to observe central black holes in high-, smaller galaxies, they had to be overmassive. However, Pacucci et al. (2023) recently found that the entire dataset of galactic systems found thus far by JWST at is so overmassive, and JWST’s sensitivity in H so deep, that their ratios cannot be explained by being extreme outliers from the local relation. They seem to be, in fact, a different statistical population of objects, with black hole masses significantly overmassive. Intriguingly, these galactic systems seem to be located close to the local and relations, suggesting that these two relationships are more fundamental than the (Maiolino et al., 2023; Juodžbalis et al., 2024). This population of sources is also X-ray weak, with several studies cross-matching samples of LRDs with X-ray catalogs and finding no or very few detections (see, e.g., Ananna et al. 2024; Yue et al. 2024a; Maiolino et al. 2024b). One possibility to solve this tension would be to assume that the IR emission is overestimating the AGN luminosity, leading to lower black hole masses. Unusually large levels of obscuration could also play a part in explaining the deficiency of X-ray emissions.
Why are these high- black holes so overmassive, and what would their subsequent cosmic evolution be? While these questions remain open and will require additional data and careful analysis, some partial answers may be already available.
For example, some studies, based on cosmological simulations (e.g., Scoggins et al. 2023) suggest that this may be the best evidence thus far available that these black holes formed at even higher redshifts ( as heavy black hole seeds (Agarwal et al., 2013). Larger samples of red and compact objects at high redshift, as well as a better understanding of the uncertainties associated with their black hole and stellar mass measurements, will be vital to confirm the very existence of this remarkable population of overmassive black holes and their broader cosmological implications.
Acknowledgments: We thank Guillermo Barro for providing the photometry of the eight sources investigated here. E.D. acknowledges support from a Dartmouth Fellowship. F.P. thanks Dale Kocevski for clarifying the differences in the selection of EROs and LRDs; our “Note on Terminology” (Sec. 2.1) is based on his insights. F.P. acknowledges support from a Clay Fellowship administered by the Smithsonian Astrophysical Observatory. R.C.H. acknowledges support from NASA through Astrophysics Data Analysis Program grant 80NSSC23K0485. This work was also supported by the Black Hole Initiative at Harvard University, which is funded by grants from the John Templeton Foundation and the Gordon and Betty Moore Foundation.